# Sharp Capacity Scaling of Spectral Optimizers in Learning Associative Memory

### 27 Mar 2026 v1

@@ Results.
Figure 6 reports the in-distribution (ID) and OOD accuracy of the three optimizers as the model dimension varies. Muon consistently outperforms SGD and AdamW across dimensions, training iterations, and both evaluation metrics. In Figures 7(a) and 7(b), we vary the batch size while fixing the number of training steps, and measure the resulting OOD accuracy and the recall accuracy of the value matrix, respectively. We again observe that Muon has the best performance; however, Muon attains near-perfect accuracy at a smaller batch size compared to SGD and AdamW, thus saturating more quickly. This discrepancy with our theoretical analysis of the critical batch size is likely due to the different task and optimizer setups, so that the information-theoretic rate is not the main bottleneck. We leave a quantitative investigation of this gap to future work.
Finally, Figure 8 demonstrates how model performance changes with the power-law exponents of the output and trigger distributions, $\alpha_{o}$ and $\alpha_{t}$, while all other distributions are kept uniform. Larger values of $\alpha$ correspond to faster decay in item frequencies and should therefore make learning more difficult. However, our discussion of signal amplification in Section 5 suggests that Muon should be more robust to this effect. Consistent with this intuition, we observe that Muon and AdamW perform similarly when $\alpha=0$, but as $\alpha$ increases, Muon remains remarkably robust whereas AdamW degrades much more sharply. This behavior also qualitatively aligns with the dependence of our one-step analysis: as $\alpha$ grows larger, Muon still retains $\Omega(d)$ capacity, but the SGD recovery rate will collapse to zero.
@@ 7 Conclusion
In this work, we study the storage capacities of Muon and SGD in a linear associative memory model with random embeddings and power-law item frequencies. We sharply characterize the one-step recovery rate of both optimizers and show that Muon recovers substantially more items than SGD. This gap also implies a much larger critical batch size for Muon, helping explain why its advantage is most visible in large-batch training. We further show that Muon’s gains are concentrated early in training, whereas over longer horizons its recovery dynamics become comparable to those of SGD. Overall, our analysis opens the door towards establishing scaling laws for spectral optimizers in realistic language modeling settings. Promising directions for future work include studying anisotropic embeddings, momentum [12], and more complex language-modeling tasks such as compositional and multi-hop reasoning [46, 57].
@@ Acknowledgements
Appendix A Proof of Theorem 3
@@
In Sections A.1-A.3, we lower bound the signal terms $\gamma_{ii}$. In Section B, we upper bound the magnitude of the interaction terms $\gamma_{ij}$ for $i\neq j$. Finally in Section A.4, we conclude the proof of Theorem 3.
@@ A.1 Fréchet derivative computations
Proposition 11 (Daleckii–Krein formula).
@@
Since ${\mathbf{T}}_{q}$ is diagonalized as ${\mathbf{T}}_{q}={\mathbf{B}}_{q}{\mathbf{S}}_{q}^{2}{\mathbf{B}}_{q}^{\top}$, we compute via the Daleckii–Krein formula (Proposition 11),
@@
Proof of Lemma 13.
Recall the definitions of ${\mathbf{G}}_{q},{\mathbf{A}}_{q},{\mathbf{S}}_{q},{\mathbf{B}}_{q},{\mathbf{Y}}_{q},{\mathbf{T}}_{q},{\mathbf{R}}_{q},a,b,\xi$ from the proof of Lemma 12. Differentiating ${\mathbf{Y}}_{q}$ twice with respect to $q$, we obtain
@@
For Eq. (11), we showed in the proof of Lemma 12 that
@@ A.2 Minibatch concentration
Lemma 14.
Define the empirical covariance matrix
@@
We first control the bulk sum. Since $|I_{k}|\leq d$, it follows from Vershynin [51, Theorem 4.6.1] that
@@
by taking $r\asymp\log d$, hence $q_{i}\lesssim\frac{\log d}{B}$. It follows from Vershynin [51, Remark 4.7.3] that
@@
Lemma 15 (concentration of tail frequencies).
@@
For $\norm{q_{>r}}_{2}$, we repeat the analysis from the proof of Lemma 14. If $B\gtrsim r^{\alpha}$, thresholding at $B^{1/\alpha}$ gives $q_{i}\lesssim p_{i}+\frac{\log d}{B}$ for items with $r<i\leq B^{1/\alpha}$, and $q_{i}\lesssim\frac{\log d}{B}$ for the $N_{\operatorname{tail}}\asymp B^{1/\alpha}$ items with $i>B^{1/\alpha}$. Combining, we obtain
@@
By Lemma 15, we have that when $B\gtrsim r^{\alpha}$,
@@ A.3 Lower bounding the signal
We now consider the stabilized sign map $h_{\lambda}(z)=\frac{z}{\sqrt{z^{2}+\lambda^{2}}}$. Recall from Lemma 12 that
@@
${\mathbf{G}}_{-i}\mathrel{\stackrel{{\scriptstyle d}}{{=}}}\frac{1}{\sqrt{d}}{\mathbf{M}}_{-i}^{1/2}{\mathbf{Z}},\quad\text{where}\quad{\mathbf{Z}}_{k\ell}\sim\mathcal{N}(0,1)\;\;\text{i.i.d.}$
It holds that [45, Eq. 2.3]
@@
Then by Lemma 14, we have $\lambda_{d/2}({\mathbf{M}})\lesssim d^{-2\alpha}(\log d)^{2}$ and so $s_{d/2}({\mathbf{G}}_{-i})\lesssim d^{-\alpha}\log d\lesssim\lambda$. We conclude from Eq. (17),
@@
Since $\phi$ is increasing by Lemma 12, we can therefore Taylor expand $\phi$ to obtain
@@ A.4 Putting things together
In Section B, we analyze the interaction terms and show that under $\mathcal{E}_{q}$ (Proposition 18),
@@
A.5 Proof of Corollary 4
Let $\hat{p}_{1}:=\hat{p}_{{\mathbf{W}}_{1}}$ be the predicted score under ${\mathbf{W}}_{1}$. By choosing $\eta\asymp(\log d)^{-4}\sqrt{d}$, we can guarantee a logit gap to
@@ B.1 Overview
We have verified $\mathcal{E}_{q}$ occurs with high probability by a judicious choice of $\lambda$ in Lemma 16, and assume this for fixed $q$ throughout the section. When either $q_{i}$ or $q_{j}\ll\lambda$, the interaction terms can be bounded by a simple operator Lipschitz concentration argument, which we provide in Section B.8. The main challenge arises when controlling the leading $r\times r$ block, where any operator norm bound fails to capture the correct scale. The analysis for these ‘large’ interactions requires a much more involved perturbative approach, and will be developed throughout Sections B.2-B.7. For the readers’ convenience, we provide a sketch of the argument here.
@@ B.2 Setup and norm estimates
Our goal is to prove Proposition 18 for the case $i,j\leq r$, which corresponds to bounding the off-diagonal entries of $\mathbf{K}$}. Set
@@
Lemma 19 (decay estimate for ${\mathbf{E}}_{u},{\mathbf{E}}_{v}$).
@@
Lemma 20 (decay estimate for $\mathbf{Z}$}).
@@
From Lemma 19 and Lemma 20, under the event $\mathcal{E}_{q}$, we can choose the proportionality constant in Eq. (24) so that $\Pr(\mathcal{E}_{\mathrm{op}})\geq 1-e^{-\Omega(r)}$. We note that the truncated series expansions in Section B.3-B.4 are valid conditional on $\mathcal{E}_{\mathrm{op}}$, however once we algebraically reduce to the appropriate quantities, we do not condition on $\mathcal{E}_{\mathrm{op}}$ for the moment computations in Section B.5-B.7.
@@
Lemma 21 (series expansion for ${\mathbf{G}}_{u},{\mathbf{G}}_{v}$).
@@
Then by Higham [17, Theorem 4.8],
@@ B.3 Block resolvent integral representation
Expanding Eq. (27) via Lemma 21, we have
@@
and ${\mathbf{R}}_{h}$ is the error term due to applying the series truncation in Lemma 21. We control truncation errors in Lemma 23 below. Next, from
@@
where the error ${\mathbf{R}}_{\delta,s}$ (here dependent on $s$) is also controlled in Lemma 23. Hence,
@@
We justify this expansion in Lemma 22 and show $\lVert{\mathbf{K}}-\tilde{}{\mathbf{K}}\rVert_{\mathrm{op}}=d^{-\omega(1)}$ in Lemma 23. Hence it suffices to bound the off-diagonal entries of $\tilde{}{\mathbf{K}}$.
@@ B.4 Truncation error bounds
Lemma 22 (Neumann series stability).
@@
Lemma 23 (truncation error bound).
@@
We have that $\lVert{\mathbf{N}}_{1}\rVert_{\mathrm{op}},\lVert{\mathbf{N}}_{2}\rVert_{\mathrm{op}}=1+o(1)$ due to Lemma 22 and Eq. (32). Moreover,
@@ B.6 Positive path correlation and graded recombination
Importantly, $\theta_{\iota}^{\mu,\nu}$ are uniformly bounded: by Lemma 24, there exists an index $m$ such that
@@
Lemma 25 (positive path correlation).
@@
Then we can bound using Lemma 25 and $|\theta_{\iota}^{\mu,\nu}|,|\theta_{\iota^{\prime}}^{\mu^{\prime},\nu^{\prime}}|\leq 1$,
@@ B.7 Graded tail bounds and hypercontractivity
For ${\mathbf{E}}_{u}$ and ${\mathbf{E}}_{v}$, recall from Lemma 19 that
@@
For $\tilde{}{\mathbf{Z}}$, we have from Lemma 20 and Eq. (16) that
@@ B.8 Lipschitz concentration for tail logits
Proposition 26 (operator Lipschitz bound).
@@
Then by Proposition 26 and $\norm{\iota(u)}_{2}\leq 2$,
@@
By Lemma 27 and concentration of Lipschitz functions of Gaussians [54, Theorem 2.26], it follows that
@@
Appendix C Proofs for SGD and Optimality
C.1 Proof of Theorem 5
@@
Finally, if $B\lesssim\sqrt{d}(\log d)^{\alpha+1}$, repeating the analysis for Lemma 15 shows that we sample at most $O(B^{1/\alpha})$ items such that $i>B^{1/\alpha}$. Aside from these items, $q_{i}=0$ and so Eq. (56) implies $\langle v_{i},v_{j}\rangle\leq d^{-1+o(1)}$, hence the same conclusion as above holds.

### 28 Apr 2026 v2

@@ Results.
Figure 7 reports the in-distribution (ID) and OOD accuracy of the three optimizers as the model dimension varies. Muon consistently outperforms SGD and AdamW across dimensions, training iterations, and both evaluation metrics. In Figures 8(a) and 8(b), we vary the batch size while fixing the number of training steps, and measure the resulting OOD accuracy and the recall accuracy of the value matrix, respectively. We again observe that Muon has the best performance; however, Muon attains near-perfect accuracy at a smaller batch size compared to SGD and AdamW, thus saturating more quickly. This discrepancy with our theoretical analysis of the critical batch size is likely due to the different task and optimizer setups, so that the information-theoretic rate is not the main bottleneck. We leave a quantitative investigation of this gap to future work.
Finally, Figure 9 examines how model performance changes with the power-law exponents of the output and trigger distributions, $\alpha_{o}$ and $\alpha_{t}$, while all other distributions are kept uniform. Larger values of $\alpha$ correspond to faster decay in item frequencies and therefore make learning more difficult. However, our discussion of signal amplification in Section 5.2 suggests that Muon should be more robust to this effect. Consistent with this intuition, we observe that Muon and AdamW perform similarly when $\alpha=0$, but as $\alpha$ increases, Muon remains remarkably robust whereas AdamW degrades much more sharply. This behavior also qualitatively aligns with our one-step analysis: as $\alpha$ grows larger, Muon still retains $\Omega(d)$ capacity, but the SGD recovery rate will collapse to zero.
@@ 7 Conclusion
In this work, we study the storage capacities of Muon, SGD, and Newton’s method in a linear associative memory model with random embeddings and power-law item frequencies. We sharply characterize the one-step recovery rate of each optimizer and show that Muon recovers substantially more items than SGD, even matching the Newton update. This gap also implies a much larger critical batch size for Muon compared to SGD, helping explain why its advantage is most visible in large-batch training. We further show that Muon’s gains are concentrated early in training, whereas over longer horizons its recovery dynamics become comparable to those of SGD. Overall, our analysis opens the door towards establishing scaling laws for spectral optimizers in realistic language modeling settings. Promising directions for future work include studying anisotropic embeddings, the role of momentum [13], and more complex language-modeling tasks such as compositional and multi-hop reasoning [47, 58].
@@ Acknowledgements
Appendix A Proof of Theorem 4.1
@@
In Sections A.1-A.3, we lower bound the signal terms $\gamma_{ii}$. In Section B, we upper bound the magnitude of the interaction terms $\gamma_{ij}$ for $i\neq j$. Finally in Section A.4, we conclude the proof of Theorem 4.1.
@@ A.1 Fréchet derivative computations
Proposition A.1 (Daleckii–Krein formula).
@@
Since ${\mathbf{T}}_{q}$ is diagonalized as ${\mathbf{T}}_{q}={\mathbf{B}}_{q}{\mathbf{S}}_{q}^{2}{\mathbf{B}}_{q}^{\top}$, we compute via the Daleckii–Krein formula (Proposition A.1),
@@
Proof of Lemma A.3.
Recall the definitions of ${\mathbf{G}}_{q},{\mathbf{A}}_{q},{\mathbf{S}}_{q},{\mathbf{B}}_{q},{\mathbf{Y}}_{q},{\mathbf{T}}_{q},{\mathbf{R}}_{q},a,b,\xi$ from the proof of Lemma A.2. Differentiating ${\mathbf{Y}}_{q}$ twice with respect to $q$, we obtain
@@
For Eq. (14), we showed in the proof of Lemma A.2 that
@@ A.2 Minibatch concentration
Lemma A.4.
Define the weighted covariance matrix
@@
We first control the bulk sum. Since $|I_{k}|\leq d$, it follows from Vershynin [52, Theorem 4.6.1] that
@@
by taking $r\asymp\log d$, hence $q_{i}\lesssim\frac{\log d}{B}$. It follows from Vershynin [52, Remark 4.7.3] that
@@
Lemma A.5 (concentration of tail frequencies).
@@
For $\norm{q_{>r}}_{2}$, we repeat the analysis from the proof of Lemma A.4. If $B\gtrsim r^{\alpha}$, thresholding at $B^{1/\alpha}$ gives $q_{i}\lesssim p_{i}+\frac{\log d}{B}$ for items with $r<i\leq B^{1/\alpha}$, and $q_{i}\lesssim\frac{\log d}{B}$ for the $N_{\operatorname{tail}}\asymp B^{1/\alpha}$ items with $i>B^{1/\alpha}$. Combining, we obtain
@@
By Lemma A.5, we have that when $B\gtrsim r^{\alpha}$,
@@ A.3 Lower bounding the signal
We now consider the stabilized sign map $h_{\lambda}(z)=\frac{z}{\sqrt{z^{2}+\lambda^{2}}}$. Recall from Lemma A.2 that
@@
${\mathbf{G}}_{-i}\mathrel{\stackrel{{\scriptstyle d}}{{=}}}\frac{1}{\sqrt{d}}{\mathbf{M}}_{-i}^{1/2}{\mathbf{Z}},\quad\text{where}\quad{\mathbf{Z}}_{k\ell}\sim\mathcal{N}(0,1)\;\;\text{i.i.d.}$ (23)
It holds that [46, Eq. 2.3]
@@
Then by Lemma A.4, we have $\lambda_{d/2}({\mathbf{M}})\lesssim d^{-2\alpha}(\log d)^{2}$ and so $s_{d/2}({\mathbf{G}}_{-i})\lesssim d^{-\alpha}\log d\lesssim\lambda$. We conclude from Eq. (20),
@@
Since $\phi$ is increasing by Lemma A.2, we can therefore Taylor expand $\phi$ to obtain
@@ A.4 Putting things together
In Section B, we analyze the interaction terms and show that under $\mathcal{E}_{q}$ (Proposition B.1),
@@
A.5 Proof of Corollary 4.2
Let $\hat{p}_{1}:=\hat{p}_{{\mathbf{W}}_{1}}$ be the predicted score under ${\mathbf{W}}_{1}$. By choosing $\eta\asymp(\log d)^{-4}\sqrt{d}$, we can guarantee a logit gap of
@@ B.1 Overview
We have verified $\mathcal{E}_{q}$ occurs with high probability by a judicious choice of $\lambda$ in Lemma A.6, and assume this for fixed $q$ throughout the section. When either $q_{i}$ or $q_{j}\ll\lambda$, the interaction terms can be bounded by a simple operator Lipschitz concentration argument, which we provide in Section B.8. The main challenge arises when controlling the leading $r\times r$ block, where any operator norm bound fails to capture the correct scale. The analysis for these ‘large’ interactions requires a much more involved perturbative approach, and will be developed throughout Sections B.2-B.7. For the readers’ convenience, we provide a sketch of the argument here.
@@ B.2 Setup and norm estimates
Our goal is to prove Proposition B.1 for the case $i,j\leq r$, which corresponds to bounding the off-diagonal entries of $\mathbf{K}$}. Set
@@
Lemma B.2 (decay estimate for ${\mathbf{E}}_{u},{\mathbf{E}}_{v}$).
@@
Lemma B.3 (decay estimate for $\mathbf{Z}$}).
@@
From Lemma B.2 and Lemma B.3, under the event $\mathcal{E}_{q}$, we can choose the proportionality constant in Eq. (28) so that $\Pr(\mathcal{E}_{\mathrm{op}})\geq 1-e^{-\Omega(r)}$. We note that the truncated series expansions in Section B.3-B.4 are valid conditional on $\mathcal{E}_{\mathrm{op}}$, however once we algebraically reduce to the appropriate quantities, we do not condition on $\mathcal{E}_{\mathrm{op}}$ for the moment computations in Section B.5-B.7.
@@
Lemma B.4 (series expansion for ${\mathbf{G}}_{u},{\mathbf{G}}_{v}$).
@@
Then by Higham [18, Theorem 4.8],
@@ B.3 Block resolvent integral representation
Expanding Eq. (31) via Lemma B.4, we have
@@
and ${\mathbf{R}}_{h}$ is the error term due to applying the series truncation in Lemma B.4. We control truncation errors in Lemma B.6 below. Next, from
@@
where the error ${\mathbf{R}}_{\delta,s}$ (here dependent on $s$) is also controlled in Lemma B.6. Hence,
@@
We justify this expansion in Lemma B.5 and show $\lVert{\mathbf{K}}-\tilde{}{\mathbf{K}}\rVert_{\mathrm{op}}=d^{-\omega(1)}$ in Lemma B.6. Hence it suffices to bound the off-diagonal entries of $\tilde{}{\mathbf{K}}$.
@@ B.4 Truncation error bounds
Lemma B.5 (Neumann series stability).
@@
Lemma B.6 (truncation error bound).
@@
We have that $\lVert{\mathbf{N}}_{1}\rVert_{\mathrm{op}},\lVert{\mathbf{N}}_{2}\rVert_{\mathrm{op}}=1+o(1)$ due to Lemma B.5 and Eq. (36). Moreover,
@@ B.6 Positive path correlation and graded recombination
Importantly, $\theta_{\iota}^{\mu,\nu}$ are uniformly bounded: by Lemma B.7, there exists an index $m$ such that
@@
Lemma B.8 (positive path correlation).
@@
Then we can bound using Lemma B.8 and $|\theta_{\iota}^{\mu,\nu}|,|\theta_{\iota^{\prime}}^{\mu^{\prime},\nu^{\prime}}|\leq 1$,
@@ B.7 Graded tail bounds and hypercontractivity
For ${\mathbf{E}}_{u}$ and ${\mathbf{E}}_{v}$, recall from Lemma B.2 that
@@
For $\tilde{}{\mathbf{Z}}$, we have from Lemma B.3 and Eq. (19) that
@@ B.8 Lipschitz concentration for tail logits
Proposition B.9 (operator Lipschitz bound).
@@
Then by Proposition B.9 and $\norm{\iota(u)}_{2}\leq 2$,
@@
By Lemma B.10 and concentration of Lipschitz functions of Gaussians [55, Theorem 2.26], it follows that
@@
Appendix C Proofs for SGD and Newton
C.1 Proof of Theorem 4.3
@@ Appendix C Proofs for SGD and Newton
Finally, if $B\lesssim\sqrt{d}(\log d)^{\alpha+1}$, repeating the analysis for Lemma A.5 shows that we sample at most $O(B^{1/\alpha})$ items such that $i>B^{1/\alpha}$. Aside from these items, $q_{i}=0$ and so Eq. (60) implies $\langle v_{i},v_{j}\rangle\leq d^{-1+o(1)}$, hence the same conclusion as above holds.